OpenStudy (anonymous):

How to factorise x^2 + 6x + 9 - y^2

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

\[(x-y+3) (x+y+3)\]

OpenStudy (anonymous):

x^2 + 6x + 9 - y^2=(x^2 + 6x + 9 )- y^2=(x^2 + 2*3*x + 9) - y^2=(x+3)^2 -y^2 =(x+3+y)(x+y-3)

OpenStudy (anonymous):

Isn't it supposed to be (x+3-y) (x-3+y)???

OpenStudy (anonymous):

(x−y+3)(x+y+3)?